A SEMIEMPIRICAL PES FOR Kr-HS

\underline{Prakashan P. Korambath}, X. T. Wu, E. F. Hayes, C. Carter and Terry. A. Miller, Dept. of Chemistry, The Ohio State University, Columbus, OH 43210.

A semiempirical potential energy surface for Kr-HS van der Waals type molecule will be presented. The nature of the PES is similar to that proposed for Ar-HO by Bowman and co-workers [1]. The parameters in the Ar-HO surface was modified using an efficient nonlinear fit program to agree with experimental measurements for Kr-HS molecule by Miller and co–workers [2]. The calculated results will be reported along with comparisons with the experimental results of Miller and co–workers. Quantitative insights into the interaction potentials for similar systems will be discussed.

1. J. M. Bowman, B. Gazdy, P. Schafer, M. C. Heaven. J. Phys. Chem 94 2226 (1990).

2. Private communication from T. A. Miller, Ohio State University.